You are here: Foswiki>Commissioning Web>Com20110702>Com20110703 (03 Jul 2011, AndrewRakich)Edit Attach

Bino collimation and pointing night 2

Observer(s): A. Rakich, K. Boutsia, O. Kuhn (LBTO)
Telescope Operator: S. Allanson (LBTO)
Telescope Support: J. Little (LBTO)
Instrument Support: none


Got binocular range balancing working for Gregorian!!! Used AGws and judicious switching on and off of the PCS/PSF automatic pointing offsets. Realized that there was no point in gathering any sort of pointing data, binocular or other, as during the summer shutdown new range balanced collimation models for all focal stations will be produced, to be implemented during telescope wake up. These will need new pointing data to be taken after they are implemented.

So, after verifying that range balancing was working, we switched to thermally cycling the primary mirrors while running active optics in Gregorian mode. Got almost half a temperature cycle.



All times are UTC unless otherwise noted.

03:10 Started trying to use AGws for binocular Gregorian range balancing. A three step process ended up giving us what we wanted:

1) Using the Lucifer and IRTC pointing models respectively for left and right, sent acquire presets and jirggered pointing into place. Then sent a guide preset with "send guide corrections" switched off in the GCS GUI. Then cancelled the preset.
2) Using the two acquisition images obtained in the guide preset in (1), do range bal.
3) Send a new guide preset to verify copointing has been achieved. this time, in the same instant the guide preset is sent, both sides tip/tilt corrections are disabled (line 35 in the pcsclient).

Very pleased to have this result after a frustrating night last night, and the bonus is that working with AGws makes this instrument independant to a large extent.

05:50 Sending active preset to circumpolar star ACT0411, GS1. This is a circumpolar star for doing thermal cycling on primaries/ active optics monitoring. Will run for rest of night.

09:33 Humidity 94%.

10:33 cloud

11:00 cloud clears, resend active preset. Had trouble acquiring left star, note added to IT 3457 for next action by Torsten. Big jump in M2 Z in expected direction given that M1 SX and DX had been cooling strongly during the last 30 minutes.

11:10 Humidity up to 96%, closing and calling it a night.

-- AndrewRakich - 03 Jul 2011
Topic revision: r3 - 03 Jul 2011, AndrewRakich
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback